OpenStudy (anonymous):

What is the distance from (−2, 4) to (0, −6)? Round your answer to the nearest hundredth

OpenStudy (irishboy123):

use Pythagoreas sqrt( (x2 - x1)^2 + (y2 - y1) ^2 )

jigglypuff314 (jigglypuff314):

Hello @abbyhenson221 and Welcome to OpenStudy! :) you can use the Distance Formula for this: \[distance formula: \sqrt{(x _{2}-x _{1})^{2} + (y _{2}-y _{1})^{2}}\]when given the coordinates \((x_1,~y_1)\) and \((x_2,~y_2)\)

jigglypuff314 (jigglypuff314):

Hint: you will get \[\sqrt{(0-(-2))^2+((-6)-4)^2}~=\]try simplifying this :)